# Post-Newtonian templates for phase evolution of spherical extreme mass ratio inspirals

### v2

@@ III.3 TaylorT2
In the circular orbit case, the TaylorT2 approximant gives $t$ and $\Phi$ as functions of $x$ which is regarded as the parametric variable. To obtain a similar expression for the spherical orbit case, first we derive the equation of $Y(x)$ as
$\frac{dY}{dx}$ $=$ $\frac{dY}{dt}\left(\frac{dx}{dt}\right)^{-1}$ (32)
@@
We can transform Eq. (33) into the form of integral equation as
@@
Here the constant term, $\tilde{Y}_{0}$, is determined so that $Y^{(n)}(x_{I})=Y_{I}$. The superscript $(n)$ in the left hand side of the above expression again means the truncation order as explained below Eq. (21). Also, note that the coefficients $\tilde{Y}_{k}$ depend on $\ln x$ in general (later, we find that the dependence of $\ln x$ appears for $k\geq 9$). By substituting Eq. (35) to $Y$ in the integral equation (34) and comparing both sides by order of $x$, we find $\tilde{Y}_{k}$ iteratively. We show the explicit formulas of $\tilde{Y}_{k}$ for $k\leq 9$ in Appendix D.
Once we obtain the expression of $Y(x)$, the equations of $t(x)$ and $\Phi(x)$ are derived with respect to $x$ as
$\frac{dt_{\textrm{T2}}^{(n)}}{dx}$ $=$ $\left(\frac{dx}{dt}\biggr{|}_{Y=Y^{(n)}(x)}\right)^{-1}=\frac{5}{32x^{9}}\frac{M^{2}}{\mu}\left[1+\frac{743}{336}x^{2}-\left(4\pi-\frac{73\tilde{Y}_{0}+40}{12}q\right)x^{3}+\cdots\right],$ (36)
$\frac{d}{dx}\Phi_{\textrm{T2}}^{(n)}$ $=$ $\frac{x^{3}}{M}\frac{dt_{\textrm{T2}}^{(n)}}{dx}.$ (37)
Since the right hand sides of Eqs. (36) and (37) are (Laurent) polynomials of $x$ and $\ln(x)$, we can integrate them with respect to $x$ analytically and obtain the TaylorT2 approximant as
@@ III.5 TaylorF1 and TaylorF2
The TaylorF1 approximant is given by solving these equations numerically after replacing $(E,L,\dot{E},\dot{L})$ in the right hand sides with their PN formulas, $(E^{(n)},L^{(n)},\dot{E}^{(n)},\dot{L}^{(n)})$. But, we do not treat TaylorF1 approximant in the rest of this paper.
By analogy from the TaylorT2 approximant in the time domain, the TaylorF2 approximant gives $t$ and $\psi$ as functions of $f$ through $x_{f}$. To this end, we first rewrite Eqs. (46)–(48) by expanding the right hand sides with respect to $x_{f}$ as
$\frac{dt_{\textrm{F2}}^{(n)}}{dx_{f}}$ $=$ $\frac{5}{32x_{f}^{9}}\frac{M^{2}}{\mu}\left[1+\frac{743}{336}x_{f}^{2}-\left(4\pi-\frac{73Y+40}{12}q\right)x_{f}^{3}+\cdots+\left(\mbox{truncated at}\ O(x_{f}^{n+1})\right)\right],$ (49)
$\frac{dY^{(n)}}{dx_{f}}$ $=$ $-q(1-Y^{2})\left[\frac{61}{24}x_{f}^{2}-\frac{13}{96}qYx_{f}^{3}-\frac{80209}{8064}x_{f}^{4}+\cdots+\left(\mbox{truncated at}\ O(x_{f}^{n+1})\right)\right].$ (50)
$\frac{d\psi_{\textrm{F2}}^{(n)}}{dx_{f}}$ $=$ $\frac{6}{M}x_{f}^{2}t_{\textrm{F2}}^{(n)}.$ (51)
Note that Eqs. (49) and (50) are equivalent to Eqs. (36) and (33), respectively, except for replacing $x$ to $x_{f}$. Hence, we can obtain $Y(x_{f})$ and $t(x_{f})$ by replacing $x$ to $x_{f}$ in Eqs. (35) and (38) as
@@
Substituting $t_{\textrm{F2}}^{(n)}$ into the right hand side of Eq. (51) and integrating with respect to $x_{f}$, we find the TaylorF2 phase as
@@ III.6 Alternative TaylorT1 and TaylorT4
$\dot{\Phi}_{\textrm{T1a}}$ $=$ $\frac{x^{3}}{M},$ (60)
where a prime ($\prime$) denotes total derivative with respect to $x$. We refer to the numerical solution of these equations as the alternative TaylorT1 (or TaylorT1a) approximant.
As in the TaylorT4 approximant, we may choose to solve Eqs. (59) and (60) after expanding the right hand sides with respect to $x$:
@@
$\dot{\Phi}_{\textrm{T4a}}$ $=$ $\frac{x^{3}}{M},$ (62)
where the expansion coefficients, $\tilde{\dot{x}}_{k}^{\textrm{T4a}}$, are given in Appendix I. We name the solution by this way the alternative TaylorT4 (or TaylorT4a) approximant.
@@ IV Results
Following Ref. [22], we consider two kinds of EMRIs, called “System1” with masses $(\mu,M)=(10,10^{5})M_{\odot}$ and “System2” with masses $(\mu,M)=(10,10^{6})M_{\odot}$. We assume that the observation of the GW starts when the GW frequency reaches $f_{I}=4.0\times 10^{-3}$ Hz (corresponding to $x_{I}\sim 0.1836$) for System1, $f_{I}=1.8\times 10^{-3}$ Hz ($x_{I}\sim 0.3031$) for System2, and lasts two years. The frequency after two-year observation reaches about $f_{\textrm{fin}}\sim 8.0\times 10^{-3}$ ($x_{\textrm{fin}}\sim 0.23$) in the case of System1 and $f_{\textrm{fin}}\sim 2.6\times 10^{-3}$ ($x_{\textrm{fin}}\sim 0.34$) in the case of System2 (strictly speaking, $f_{\textrm{fin}}$ (and $x_{\textrm{fin}}$) depends on the values of $(q,Y_{I})$). System1 represents the early stage of EMRIs, while System2 does the late stage.
IV.1 Convergence of PN formulas for $E$ and $L$
To investigate the convergence of the PN formulas of the specific energy and angular momentum, Eqs. (20) and (21), we introduce the difference between the PN formulas with two successive orders of $x$ as
Since $\Delta E^{(n)}$ ($\Delta L^{(n)}$) corresponds to the $(n/2)$-th PN correction in $E$ ($L$), it decreases as $n$ increases if the PN formula converges. In Figs. 1 and 2, we show $\Delta E^{(n)}$ and $\Delta L^{(n)}$ as functions of $x$ for several sets of $(q,Y)$. In these plots, $\Delta E^{(n)}$ and $\Delta L^{(n)}$ are normalized by the leading (Newtonian) order terms,
$E_{N}\equiv 1-\frac{1}{2}x^{2},\quad L_{N}\equiv\frac{Y}{x}.$
We find a weak dependence on $q$ in Figs. 1 and 2: the PN convergence gets slightly worse when $q$ increases. On the other hand, we find no dependence on $Y$ in these plots at least with one’s eyes.
@@ IV.1 Convergence of PN formulas for $E$ and $L$
Figure 1: Contribution of each correction term in the PN expansion of the specific energy. We plot $\Delta E^{(n)}$ normalized by the Newtonian energy, $E_{N}$, as functions of $x$ for several sets of $(q,Y)$. To make the plots easier to read, we show only the 2PN ($n=4$), 4PN ($n=8$), 6PN ($n=12$), 8PN ($n=16$), 10PN ($n=20$) and 12PN ($n=24$) corrections in this figure.
IV.2 Convergence of PN formulas for $\dot{E}$ and $\dot{L}$
In a similar manner to Eq. (63), we check the convergence of the PN formulas of $\dot{E}$ and $\dot{L}$ by introducing the difference
@@ IV.2 Convergence of PN formulas for $\dot{E}$ and $\dot{L}$
In Figs. 3 and 4, we show $\Delta\dot{E}^{(n)}$ and $\Delta\dot{L}^{(n)}$ as functions of $x$ for several sets of $(q,Y)$. We find some spike bottoms which are caused by the logarithmic terms appearing in the PN formulas. Unlike $E$ and $L$, it is difficult to find any tendency in the convergence of $\dot{E}$ and $\dot{L}$ depending on $q$ and $Y$. The convergence is not uniform comparing to $E^{(n)}$ and $L^{(n)}$, but $\Delta\dot{E}^{(n)}$ and $\Delta\dot{L}^{(n)}$ tend to decrease when the PN order increases. Therefore, we expect to obtain more accurate approximants at higher PN order. It is noted that the convergence of $\Delta\dot{E}^{(n)}$ and $\Delta\dot{L}^{(n)}$ may be improved by using resummation technique [22], although we do not deal with this topic in this paper.
Figure 3: Contribution of each correction term in the PN expansion of $\dot{E}$. We plot $\Delta\dot{E}^{(n)}$ normalized by the leading term, $\dot{E}^{(0)}$, as functions of $x$ for several sets of $(q,Y)$. In the same reason as Fig. 1, we show only the 2PN, 4PN, 6PN, 8PN, 10PN and 12PN corrections in this figure.
Figure 4: Contribution of each correction term in the PN expansion of $\dot{L}$. We plot $\Delta\dot{L}^{(n)}$ normalized by the leading term, $\dot{L}^{(0)}$, as functions of $x$ for several sets of $(q,Y)$. In the same reason as Fig. 1, we show only the 2PN, 4PN, 6PN, 8PN, 10PN and 12PN corrections in this figure.
IV.3 Comparison of the time-domain Taylor approximants
In a similar manner to Eq. (63), we introduce the difference in the phase between two PN formulas with different order of $x$ for each TaylorT approximant:
@@ IV.3 Comparison of the time-domain Taylor approximants
Figures 59 show $\Delta\Phi_{X}^{(n)}$ of each approximant as functions of time (in unit of month) for several sets of $(q,Y_{I})$ in System1. We find in all cases of $(q,Y_{I})$ that the TaylorT1 shows best convergence. The convergence of the TaylorT1a (TaylorT4a) is almost same as the TaylorT1 (TaylorT4). This suggests that the PN formula of $Y$, Eq. (35), converges well. The TaylorT2 approximant also shows good performance though the convergence is slightly slower than the TaylorT1. The convergence of the TaylorT3 and TaylorT4 is not good comparing to TaylorT1 and TaylorT2. Especially, the TaylorT3 shows poor convergence comparing to the other approximants. This trend is consistent with the result for circular orbits in Schwarzschild case [29]. The convergence of TaylorT3 and TaylorT4 gets slightly better when $q$ and $Y_{I}$ are smaller.
In all plots of Figs. 59, we show the line with $\Delta\Phi_{X}^{(n)}=0.1$ as a reference, which corresponds to the statistical error of the phase estimated by $\langle\Delta\Phi\rangle\sim 1/\rho$ with $\rho\simeq 20$ ($\rho$ is the signal-to-noise ratio (SNR)) [33, 34, 35]. As for the TaylorT1 approximant, the 8PN correction of the phase is less than $O(0.1)$ for two years since the GW observation starts for all cases of System1, while the 9th or higher PN corrections in the TaylorT2 are required to suppress the dephasing to $O(0.1)$. Roughly speaking, the following two higher order corrections are needed to obtain the comparable accuracy by using the TaylorT2 approximant instead of the TaylorT1.
We show the results for System2 in Figs. 1012. In the case of System2, even the 12PN correction, the highest order term we have in hand currently, cannot keep the dephasing less than $O(0.1)$ for two years. This is expected because System2 corresponds to the late stage of a inspiral just before plunging into the central black hole, where the PN convergence gets worse. The tendency of the slow convergence is more pronounced for the TaylorT3 and TaylorT4 approximants. Figures 1012 do not include the results for the TaylorT3 approximant because the values are far from the other approximants in the case of System2 and the comparison does not make sense.) The reason has been discussed for circular orbit in Schwarzschild case [29]: the slow PN convergence in the late inspiral is attributed to the pole of $(dE/dx)^{-1}$ at the last stable orbit (LSO). In a similar manner, the inverse of the Jacobian, $G^{A}_{B}$, appearing in Eqs. (17) and (18), is expected to have a pole at the LSO, and the PN convergence is slow near the pole. Some techniques of resummation or factorization may be required to improve the slow convergence around the LSO [36, 37].
@@
Figure 5: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.1$ and $Y_{I}=0.9$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=10^{-1}$.
@@
Figure 6: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.5$ and $Y_{I}=0.9$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=10^{-1}$.
@@
Figure 7: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.9$ and $Y_{I}=0.9$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=10^{-1}$.
Figure 8: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.9$ and $Y_{I}=0.5$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=10^{-1}$.
Figure 9: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.9$ and $Y_{I}=0.1$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=10^{-1}$.
@@
Figure 10: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System2 with $q=0.9$ and $Y_{I}=0.9$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=10^{-1}$. The plot of the TaylorT3 approximants is not shown because most of them are out of the common range with the other approximants, $10^{-3}\leq\Delta\Phi^{(n)}\leq 10^{4}$, and the comparison does not make sense.
@@
Figure 11: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System2 with $q=0.9$ and $Y_{I}=0.5$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=10^{-1}$.
@@
Figure 12: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System2 with $q=0.9$ and $Y_{I}=0.1$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=10^{-1}$.
IV.4 Performance of the frequency-domain Taylor approximants

### v3

@@ III.3 TaylorT2
In the circular orbit case, the TaylorT2 approximant gives $t$ and $\Phi$ as functions of $x$ which is regarded as the parametric variable. To obtain a similar expression for the spherical orbit case, first we derive the equation of $Y(x)$ in the PN form as
$\frac{dY}{dx}$ $=$ $\frac{\dot{Y}}{\dot{x}}$ (34)
@@
We can transform Eq. (35) into an integral form as
@@
Here the constant term, $\tilde{Y}_{0}$, is determined so that $Y^{(n)}(x_{I})=Y_{I}$. The superscript $(n)$ in the left-hand side of the above expression again means the truncation order as explained below Eq. (21). Also, note that the coefficients $\tilde{Y}_{k}$ depend on $\ln x$ in general (later, we find that the dependence of $\ln x$ appears for $k\geq 9$). By substituting Eq. (37) to $Y$ in the integral equation (36) and comparing both sides by order of $x$, we find $\tilde{Y}_{k}$ iteratively. We show the explicit formulas of $\tilde{Y}_{k}$ for $k\leq 9$ in Appendix D.
Once we obtain the expression of $Y(x)$, the set of equations for $t(x)$ and $\Phi(x)$ are derived as
$\frac{dt}{dx}$ $=$ $\left(\dot{x}\bigr{|}_{Y=Y(x)}\right)^{-1},$ (38)
$\frac{d\Phi}{dx}$ $=$ $\frac{x^{3}}{M}\frac{dt}{dx}.$ (39)
Substituting Eqs. (29) and (37) into the right-hand side of Eq. (38) and performing the PN expansion up to $O(x^{n})$ relative to the leading order, we obtain
$\frac{dt}{dx}$ $=$ $\frac{5}{32x^{9}}\frac{M^{2}}{\mu}\left[1+\frac{743}{336}x^{2}-\left(4\pi-\frac{73\tilde{Y}_{0}+40}{12}q\right)x^{3}+\cdots+\left(\mbox{truncated at}\ O(x^{n+1})\right)\right],$ (40)
$\frac{d\Phi}{dx}$ $=$ $\frac{x^{3}}{M}\frac{dt}{dx}.$ (41)
Since the right-hand sides of Eqs. (40) and (41) are (Laurent) polynomials of $x$ and $\ln(x)$, we can integrate them with respect to $x$ analytically and obtain the TaylorT2 approximant, which we represent by $t_{\textrm{T2}}^{(n)}$ and $\Phi_{\textrm{T2}}^{(n)}$, as
@@ III.5 TaylorF1 and TaylorF2
The TaylorF1 approximant is given by solving these equations numerically after replacing $(E,L,\dot{E},\dot{L})$ in the right-hand sides with their PN formulas, $(E^{(n)},L^{(n)},\dot{E}^{(n)},\dot{L}^{(n)})$. But we do not treat the TaylorF1 approximant in the rest of this paper.
By analogy from the TaylorT2 approximant in the time domain, the TaylorF2 approximant gives $t$ and $\psi$ as functions of $f$ through $x_{f}$. To this end, we first rewrite Eqs. (50)–(52) by expanding the right-hand sides with respect to $x_{f}$ as
$\frac{dt}{dx_{f}}$ $=$ $\frac{5}{32x_{f}^{9}}\frac{M^{2}}{\mu}\left[1+\frac{743}{336}x_{f}^{2}-\left(4\pi-\frac{73Y+40}{12}q\right)x_{f}^{3}+\cdots+\left(\mbox{truncated at}\ O(x_{f}^{n+1})\right)\right],$ (53)
$\frac{dY}{dx_{f}}$ $=$ $-q(1-Y^{2})\left[\frac{61}{24}x_{f}^{2}-\frac{13}{96}qYx_{f}^{3}-\frac{80209}{8064}x_{f}^{4}+\cdots+\left(\mbox{truncated at}\ O(x_{f}^{n+1})\right)\right].$ (54)
$\frac{d\psi}{dx_{f}}$ $=$ $\frac{6}{M}x_{f}^{2}t.$ (55)
Note that Eqs. (53) and (54) are equivalent to Eqs. (40) and (35), respectively, except for replacing $x$ to $x_{f}$. Hence, we can obtain the analytic solutions of Eqs. (53) and (54), represented by $Y^{(n)}(x_{f})$ and $t_{\textrm{F2}}^{(n)}(x_{f})$, by replacing $x$ to $x_{f}$ in Eqs. (37) and (42) as
@@
Substituting $t_{\textrm{F2}}^{(n)}$ into the right-hand side of Eq. (55) and integrating with respect to $x_{f}$, we find the TaylorF2 phase as
@@ III.6 Alternative TaylorT1 and TaylorT4
$\dot{\Phi}$ $=$ $\frac{x^{3}}{M},$ (64)
where a prime ($\prime$) denotes total derivative with respect to $x$. We refer to the numerical solution of these equations as the alternative TaylorT1 (or TaylorT1a) approximant and represent the solution with respect to $\Phi$ by $\Phi_{\textrm{T1a}}^{(n)}$.
As in the TaylorT4 approximant, we may choose to solve Eqs. (63) and (64) after expanding the right-hand sides of Eq. (63) with respect to $x$:
@@
$\dot{\Phi}$ $=$ $\frac{x^{3}}{M},$ (66)
where the expansion coefficients, $\tilde{\dot{x}}_{k}^{\textrm{T4a}}$, are given in Appendix I. We name the solution in this way the alternative TaylorT4 (or TaylorT4a) approximant and represent the solution with respect to $\Phi$ by $\Phi_{\textrm{T4a}}^{(n)}$.
@@ IV Results
Following Ref. [22], we consider two kinds of EMRIs, called “System1” with masses $(\mu,M)=(10,10^{5})M_{\odot}$ and “System2” with masses $(\mu,M)=(10,10^{6})M_{\odot}$. We assume that the observation of the GW starts when the GW frequency reaches $f_{I}=4.0\times 10^{-3}$ Hz (corresponding to $x_{I}\sim 0.184$) for System1, $f_{I}=1.8\times 10^{-3}$ Hz ($x_{I}\sim 0.303$) for System2, and lasts two years. The orbital parameters and the corresponding frequency of the GW after two-year observation are shown in Table 1 for System1 and Table 2 for System2. The normalized radius in System1 is relatively large ($r/M\geq 18$) compared to System2 ($r/M\leq 12$). This means that System1 is in the early stage of the inspiral, while System2 is in the late stage.
Table 1: Values of orbital parameters at start and end of observation for System1. $x_{I}$, $r_{I}$, $Y_{I}$ are the values of $x$, $r$, $Y$ at start of observation, while $x_{\textrm{fin}}$, $r_{\textrm{fin}}$, $Y_{\textrm{fin}}$ after two-year observation, which are evaluated by solving the TaylorT1 equations, Eqs. (26)–(28), at the 12PN order. The corresponding frequency to $x_{\textrm{fin}}$, $f_{\textrm{fin}}$, is also shown in this table.
$q$ $x_{I}$ $r_{I}/M$ $Y_{I}$ $x_{\textrm{fin}}$ $r_{\textrm{fin}}/M$ $Y_{\textrm{fin}}$ $f_{\textrm{fin}}$ [Hz]
$0.10$ $0.184$ $29.7$ $0.90$ $0.234$ $18.2$ $0.900$ $8.32\times 10^{-3}$
$0.10$ $0.184$ $29.7$ $0.50$ $0.235$ $18.1$ $0.500$ $8.37\times 10^{-3}$
$0.10$ $0.184$ $29.7$ $0.10$ $0.235$ $18.1$ $0.100$ $8.43\times 10^{-3}$
$0.50$ $0.184$ $29.6$ $0.90$ $0.229$ $19.0$ $0.900$ $7.75\times 10^{-3}$
$0.50$ $0.184$ $29.7$ $0.50$ $0.231$ $18.8$ $0.498$ $7.93\times 10^{-3}$
$0.50$ $0.184$ $29.8$ $0.10$ $0.233$ $18.6$ $0.098$ $8.14\times 10^{-3}$
$0.90$ $0.184$ $29.6$ $0.90$ $0.225$ $19.6$ $0.899$ $7.37\times 10^{-3}$
$0.90$ $0.184$ $29.7$ $0.50$ $0.227$ $19.4$ $0.497$ $7.58\times 10^{-3}$
$0.90$ $0.184$ $29.8$ $0.10$ $0.230$ $19.1$ $0.096$ $7.88\times 10^{-3}$
Table 2: Values of orbital parameters at start and end of observation for System2. This table is corresponding to Table 1 for System2.
$q$ $x_{I}$ $r_{I}/M$ $Y_{I}$ $x_{\textrm{fin}}$ $r_{\textrm{fin}}/M$ $Y_{\textrm{fin}}$ $f_{\textrm{fin}}$ [Hz]
$0.10$ $0.303$ $10.9$ $0.90$ $0.344$ $8.45$ $0.900$ $2.62\times 10^{-3}$
$0.10$ $0.303$ $10.9$ $0.50$ $0.346$ $8.38$ $0.500$ $2.67\times 10^{-3}$
$0.10$ $0.303$ $10.9$ $0.10$ $0.348$ $8.30$ $0.099$ $2.72\times 10^{-3}$
$0.50$ $0.303$ $10.8$ $0.90$ $0.331$ $9.05$ $0.900$ $2.34\times 10^{-3}$
$0.50$ $0.303$ $10.9$ $0.50$ $0.334$ $8.98$ $0.498$ $2.41\times 10^{-3}$
$0.50$ $0.303$ $11.0$ $0.10$ $0.340$ $8.83$ $0.098$ $2.53\times 10^{-3}$
$0.90$ $0.303$ $10.7$ $0.90$ $0.325$ $9.31$ $0.899$ $2.22\times 10^{-3}$
$0.90$ $0.303$ $10.9$ $0.50$ $0.328$ $9.32$ $0.497$ $2.28\times 10^{-3}$
$0.90$ $0.303$ $11.1$ $0.10$ $0.334$ $9.23$ $0.096$ $2.40\times 10^{-3}$
IV.1 Convergence of PN formulas for $\dot{E}$ and $\dot{L}$
@@
Figure 1: Contribution of each correction term in the PN expansion of $\dot{E}$. We plot $\Delta\dot{E}^{(n)}$ normalized by the leading term, $\dot{E}^{(0)}$, as functions of $x$ for several sets of $(q,Y)$. To make the plots easier to read, we show only the 2PN ($n=4$), 4PN ($n=8$), 6PN ($n=12$), 8PN ($n=16$), 10PN ($n=20$) and 12PN ($n=24$) corrections in this figure. Some spike bottoms are caused by the logarithmic terms appearing in the PN formulas. We can see that the range of convergence extends to larger $x$ by using higher PN order approximations.
Figure 2: Contribution of each correction term in the PN expansion of $\dot{L}$. We plot $\Delta\dot{L}^{(n)}$ normalized by the leading term, $\dot{L}^{(0)}$, as functions of $x$ for several sets of $(q,Y)$. In the same reason as Fig. 1, we show only the 2PN, 4PN, 6PN, 8PN, 10PN and 12PN corrections in this figure. Some spike bottoms are caused by the logarithmic terms appearing in the PN formulas. We can see that the range of convergence extends to larger $x$ by using higher PN order approximations.
To investigate the convergence of the PN formulas of $\dot{E}$ and $\dot{L}$, we introduce the difference between the PN formulas with two successive orders of $x$ as
@@
Since $\Delta\dot{E}^{(n)}$ ($\Delta\dot{L}^{(n)}$) corresponds to the $(n/2)$-th PN correction in $\dot{E}$ ($\dot{L}$), it decreases as $n$ increases if the PN formula converges. (See also Appendix J for the convergence of the PN formulas of ${E}$ and ${L}$.)
In Figs. 1 and 2, we show $\Delta\dot{E}^{(n)}$ and $\Delta\dot{L}^{(n)}$ as functions of $x$ for several sets of $(q,Y)$. We find some spike bottoms which are caused by the logarithmic terms appearing in the PN formulas. Unlike $E$ and $L$ presented in Appendix J, it is difficult to find any tendency in the convergence of $\dot{E}$ and $\dot{L}$ depending on $q$ and $Y$. The convergence is not uniform comparing to $E^{(n)}$ and $L^{(n)}$, but $\Delta\dot{E}^{(n)}$ and $\Delta\dot{L}^{(n)}$ tend to decrease when the PN order increases. Therefore, we expect to obtain more accurate approximants at higher PN order. The convergence of $\Delta\dot{E}^{(n)}$ and $\Delta\dot{L}^{(n)}$ might be improved by using resummation techniques [22]. This is left to future work.
IV.2 Comparison of the time-domain Taylor approximants
Figure 3: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.1$ and $Y_{I}=0.9$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=0.05$, which corresponds to the statistical error of the phase with the SNR of $\rho\simeq 20$.
Figure 4: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.5$ and $Y_{I}=0.9$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=0.05$, which corresponds to the statistical error of the phase with the SNR of $\rho\simeq 20$.
In a similar manner to Eq. (67), we introduce the difference in the phase between two PN formulas with different orders of $x$ for each TaylorT approximant:
@@ IV.2 Comparison of the time-domain Taylor approximants
Figure 5: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.9$ and $Y_{I}=0.9$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=0.05$, which corresponds to the statistical error of the phase with the SNR of $\rho\simeq 20$.
@@
Figure 6: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.9$ and $Y_{I}=0.5$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=0.05$, which corresponds to the statistical error of the phase with the SNR of $\rho\simeq 20$.
@@
Figure 7: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System1 with $q=0.9$ and $Y_{I}=0.1$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=0.05$, which corresponds to the statistical error of the phase with the SNR of $\rho\simeq 20$.
Figures 37 show $\Delta\Phi_{X}^{(n)}$ of each approximant as functions of time (in units of month) for several sets of $(q,Y_{I})$ in System1. We find in all cases of $(q,Y_{I})$ that the TaylorT1 shows best convergence. The convergence of the TaylorT1a (TaylorT4a) is almost the same as the TaylorT1 (TaylorT4). This suggests that the PN formula of $Y$, Eq. (37), converges well. The TaylorT2 approximant also shows good performance though the convergence is slightly slower than the TaylorT1. The convergence of the TaylorT3 and TaylorT4 is not good comparing to TaylorT1 and TaylorT2. Especially, the TaylorT3 shows poor convergence comparing to the other approximants. This trend is consistent with the result for circular orbits in Schwarzschild spacetime [38]  [^7] For example, Ref. [39] has shown the best performance of the TaylorT4 in the TaylorT models for comparable-mass binary black holes (see also Ref. [40] for ineffectiveness of resummation techniques in the PN expansion). Thus, we expected that the performance might vary depending on the situation being considered.. The convergence of TaylorT3 and TaylorT4 gets slightly better when $q$ and $Y_{I}$ are smaller.
In all plots of Figs. 37, we show the line with $\Delta\Phi_{X}^{(n)}=0.05$ as a reference, which corresponds to the statistical error of the phase estimated by $\langle\Delta\Phi\rangle\sim 1/\rho$ with $\rho\simeq 20$  [^8] Assuming $\rho=20$, we obtain $\langle\Delta\Phi\rangle\sim 1/\rho=0.05$. ($\rho$ is the signal-to-noise ratio (SNR)) [41, 42, 43]. As for the TaylorT1 approximant, the 8PN correction of the phase is less than $O(0.1)$ for two years since the GW observation starts for all cases of System1. On the other hand, the 9th or higher PN corrections in the TaylorT2 are required to suppress the dephasing to $O(0.1)$. Roughly speaking, the following two higher order corrections are needed to obtain the comparable accuracy by using the TaylorT2 approximant instead of the TaylorT1.
@@
Figure 8: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System2 with $q=0.9$ and $Y_{I}=0.9$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=0.05$, which corresponds to the statistical error of the phase with the SNR of $\rho\simeq 20$. The plot of the TaylorT3 approximants is not shown because most of them are out of the common range with the other approximants, $10^{-3}\leq\Delta\Phi^{(n)}\leq 10^{4}$, and the comparison does not make sense.
@@
Figure 9: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System2 with $q=0.9$ and $Y_{I}=0.5$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=0.05$, which corresponds to the statistical error of the phase with the SNR of $\rho\simeq 20$. The plot of the TaylorT3 approximants is not shown because most of them are out of the common range with the other approximants, $10^{-3}\leq\Delta\Phi^{(n)}\leq 10^{4}$, and the comparison does not make sense.
@@
Figure 10: Difference of the phase evolution, $\Delta\Phi_{X}^{(n)}$, for several TaylorT approximants in System2 with $q=0.9$ and $Y_{I}=0.1$. The horizontal dotted line denotes $\Delta\Phi_{X}^{(n)}=0.05$, which corresponds to the statistical error of the phase with the SNR of $\rho\simeq 20$. The plot of the TaylorT3 approximants is not shown because most of them are out of the common range with the other approximants, $10^{-3}\leq\Delta\Phi^{(n)}\leq 10^{4}$, and the comparison does not make sense.
We show the results for System2 in Figs. 810. In the case of System2, even the 12PN correction, the highest order term we have in hand currently, cannot keep the dephasing less than $O(0.1)$ for two years. This is expected because System2 corresponds to the late stage of an inspiral just before plunging into the central black hole, where the PN convergence gets worse. The tendency of the slow convergence is more pronounced for the TaylorT3 and TaylorT4 approximants. Figures 810 do not include the results for the TaylorT3 approximant because the values are far from the other approximants in the case of System2 and the comparison does not make sense. The reason has been discussed for circular orbit in Schwarzschild spacetime [38]: the slow PN convergence in the late inspiral is attributed to the pole of $(dE/dx)^{-1}$ at the last stable orbit (LSO). In a similar manner, the inverse of the Jacobian, $G^{A}_{B}$, appearing in Eqs. (17) and (18), is expected to have a pole at the LSO, and the PN convergence is slow near the pole. Some techniques of resummation or factorization may be required to improve the slow convergence around the LSO [44, 45].
IV.3 Performance of the frequency-domain Taylor approximants